Your body is thrown vertically up with the velocity 98 m/s much time it will rise? (g=9.8 m/s2

Grade:9

1 Answers

Arun
6 years ago
Dear student
 

Initial velocity (u) = 98 m/s    Final velocity (v) = 0 m/s

Gravity in toward down = + 9.8 m/s Gravity in toward up = - 9.8 m/s2

(i) We know, V2 - u2 = 2gs

              or,  (o)2- (98)2 = -2 X 9.8 X S

              or,  s =  (98)2/ 2 X 9.8 = 490 m

Maximum height = 490 m

 

Regards

Arun (askIITians forum expert)